z2(p2 + q2 + 1) = k2



1
Expert's answer
2021-07-21T08:50:04-0400
"f(x, y, z, p, q)=z^2(p^2+q^2+1)-k^2=0"

"f_p=2pz^2, f_q=2qz^2,"


"f_x=0, f_y=0, f_z=2z(p^2+q^2+1)"

Charpit's Method


"\dfrac{dx}{f_p}=\dfrac{dy}{f_q}=\dfrac{dz}{pf_p+qf_q}=\dfrac{dp}{-(f_x+pf_z)}=\dfrac{dq}{-(f_y+qf_z)}"

Then


"\dfrac{dp}{-(0+2pz(p^2+q^2+1))}=\dfrac{dq}{-(0+2qz(p^2+q^2+1))}"

"\dfrac{dp}{p}=\dfrac{dq}{q}"

"\int\dfrac{dp}{p}=\int\dfrac{dq}{q}"

"\ln|p|=\ln|q|+\ln a"

"p=qa"

Substitute


"z^2(a^2q^2+q^2+1)=k^2"


"q^2=\dfrac{k^2-z^2}{z^2(a^2+1)}"

"q=\sqrt{\dfrac{k^2-z^2}{z^2(a^2+1)}}"

"p=a\sqrt{\dfrac{k^2-z^2}{z^2(a^2+1)}}"

"dz=a\sqrt{\dfrac{k^2-z^2}{z^2(a^2+1)}}dx+\sqrt{\dfrac{k^2-z^2}{z^2(a^2+1)}}dy"

"z\sqrt{\dfrac{a^2+1}{k^2-z^2}}dz=adx+dy"

Integrate


"-\sqrt{a^2+1}\sqrt{k^2-z^2}=ax+y+b"

"\sqrt{a^2+1}\sqrt{k^2-z^2}+ax+y+b=0"